OpenStudy (anonymous):

Prove: (secx+cscx)cotx=1

OpenStudy (fibonaccichick666):

So what do you think you should do first?

OpenStudy (anonymous):

I did convert secx into 1/cosx and cscx into 1/sinx

OpenStudy (anonymous):

and cotx into cosx/sinx

OpenStudy (fibonaccichick666):

good start. So we want to get rid of those parentheses now...

OpenStudy (anonymous):

it's going to turn into 1/sinx + cosx/sin^2x I have no idea how it's going to get to 1

OpenStudy (fibonaccichick666):

trust me?

OpenStudy (fibonaccichick666):

Once you do that, get a common denominator

OpenStudy (anonymous):

I dont know what the next step should be...

OpenStudy (anonymous):

cosx/sinx+sin^2x

OpenStudy (fibonaccichick666):

wait, that fraction isn't right look at it and use this \frac{}{}\) put your numerator and denominator in the brackets and just add a \( in front

OpenStudy (anonymous):

actually it turns to (1+cosx)/(sinx+sin^2x) opps..

OpenStudy (fibonaccichick666):

no... Idk how you got that. The common denom is just \(sin^2x\)

OpenStudy (anonymous):

in the beginning cotx secx + cotx cscx =(cosx/sinx)*(1/cosx) + (cosx/sinx) (1/sinx)= 1/sinx + (cosx/sin^2x)

OpenStudy (fibonaccichick666):

yea, now get a common denominator

OpenStudy (anonymous):

that'll simplify to 1+cosx/sinx+sin^2x right?

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

How are you doing that?

OpenStudy (anonymous):

1/sinx + (cosx/sin^2x)= (1+cosx)/(sinx+sin^2x)

OpenStudy (fibonaccichick666):

no not at all ok I see what you did now

OpenStudy (fibonaccichick666):

so rules for fractions

OpenStudy (fibonaccichick666):

let a,b,c,d be numbers if we have \[\frac{a}{b}+\frac{c}{d}\not= \frac{a+b}{c+d}\]

OpenStudy (anonymous):

I did a + c/b + d

OpenStudy (fibonaccichick666):

Still not how we do it. We actually have to find the same denominator so we do this: \[\frac{a}{b}+\frac{c}{d}=(\frac{a}{b} \times \frac{d}{d}) +(\frac{c}{d} \times \frac{b}{b})= \frac{ad+cb}{bd}\]

OpenStudy (fibonaccichick666):

this is called finding the least common denominator. Same as if you had 1/3+1/4

OpenStudy (fibonaccichick666):

But yea, oops alphabetical moment on the wrong way

OpenStudy (anonymous):

so it's (sin^2x+cosx sinx)/ sin^3x

OpenStudy (fibonaccichick666):

no, well yes, but reduce

OpenStudy (anonymous):

I have no clue what to do next then....

OpenStudy (fibonaccichick666):

reduce

OpenStudy (anonymous):

how?..

OpenStudy (fibonaccichick666):

what term is common to each element?

OpenStudy (fibonaccichick666):

said that backwards: Which element is common to every term?

OpenStudy (anonymous):

sinx

OpenStudy (fibonaccichick666):

good so eliminate or in dr who terms EXTERMINATE!!!

OpenStudy (anonymous):

so it's sinx(sinx+cosx/sin^2x)

OpenStudy (fibonaccichick666):

yup now here is where we find magic

OpenStudy (anonymous):

I have no clue what the "magic" is now...

OpenStudy (fibonaccichick666):

wait no read that wrong

OpenStudy (fibonaccichick666):

there shouldnt be a sinx in front

OpenStudy (fibonaccichick666):

they cancel with the bottom and =1

OpenStudy (anonymous):

wait what do you mean?

OpenStudy (fibonaccichick666):

so x/x=1 y/y=1 sinx/sinx=1

OpenStudy (fibonaccichick666):

follow?

OpenStudy (anonymous):

yea...

OpenStudy (fibonaccichick666):

so our fraction becomes \[\frac{sinx(sinx+cosx)}{sinx(sin^2x)}\]

OpenStudy (fibonaccichick666):

cancels and voila just \frac{(sinx+cosx)}{(sin^2x)}

OpenStudy (fibonaccichick666):

\[\frac{(sinx+cosx)}{(sin^2x)}\]

OpenStudy (anonymous):

yea...then what's next...

OpenStudy (fibonaccichick666):

do you see the math there?

OpenStudy (anonymous):

yep

OpenStudy (fibonaccichick666):

like that we just did nothing new

OpenStudy (fibonaccichick666):

k

OpenStudy (fibonaccichick666):

meh h/o let me actually do this real quick It didn't simplify how I wanted it to in my head

OpenStudy (fibonaccichick666):

are you sure you copied the problem correctly?

OpenStudy (fibonaccichick666):

can you double check for me please?

OpenStudy (fibonaccichick666):

I think the sec and csc might have been squared

OpenStudy (anonymous):

nope they werent

OpenStudy (fibonaccichick666):

hmmm...

OpenStudy (fibonaccichick666):

tricky tricky

OpenStudy (fibonaccichick666):

This is not true

OpenStudy (anonymous):

teacher said it was a level 1 problem; I don't believe him...

OpenStudy (fibonaccichick666):

This actually isn't plausible you effectively say that : \[\frac{(sinx+cosx)}{(sin^2x)}=1\]

OpenStudy (fibonaccichick666):

which cannot be true

OpenStudy (fibonaccichick666):

for every x

OpenStudy (fibonaccichick666):

only if sinx=1 does this work

OpenStudy (fibonaccichick666):

so hmm...

OpenStudy (fibonaccichick666):

they said it was solvable?

OpenStudy (anonymous):

yea, I was supposed to prove it but, it's complicated..

OpenStudy (fibonaccichick666):

it's impossible to prove, just test it from the get go. use any x other than pi/2. it will not =1

OpenStudy (fibonaccichick666):

i mean we could try some tricks... let's see

OpenStudy (fibonaccichick666):

\[ \frac{\sqrt{sin^2x+cos^2x+2sinx cosx}}{sin^x}\]

OpenStudy (fibonaccichick666):

oops\[\frac{\sqrt{sin^2x+cos^2x+2sinx cosx}}{sin^2x}\]

OpenStudy (fibonaccichick666):

\[=\frac{\sqrt{2sinxcosx}}{sin^2x}\]

OpenStudy (fibonaccichick666):

\[==\sqrt{\frac{2sinxcosx}}{sin^4x}}

OpenStudy (fibonaccichick666):

\[=\sqrt{\frac{2sinxcosx}{sin^4x}}\]

OpenStudy (fibonaccichick666):

\[=\sqrt{\frac{2cosx}{sin^2xsinx}}\]

OpenStudy (fibonaccichick666):

\[=sinx\sqrt{2cotx}\] Which obviously doesn not =1

OpenStudy (anonymous):

my head is about to explode, it's 11pm and that's only problem #2 and I have 18 more questions to go.

OpenStudy (anonymous):

I've tried starting with sinx/cosx(1/sinx+cosx)

OpenStudy (fibonaccichick666):

It's physically impossible

OpenStudy (fibonaccichick666):

just provide a counter ex from the start. Use like pi/4 and show it doesnt work

OpenStudy (fibonaccichick666):

I mean that should show it. \[\frac{2}{\sqrt{2}}\times 1 \not= 1\]

OpenStudy (fibonaccichick666):

and tell them they screwed up

OpenStudy (fibonaccichick666):

go to the next question and realize your teacher is a jerk or a nitwit. And finish them. Don't sweat this one. Prove it wrong then move on

OpenStudy (anonymous):

lol, my teacher's already a jerk

OpenStudy (fibonaccichick666):

yea, when they said this was a level 1 proof, they must have meant a level one to disprove

OpenStudy (fibonaccichick666):

probably just did it to give themselves an ego boost

OpenStudy (anonymous):

if I have a sinx/sin^2x does it become 1/sinx?

OpenStudy (fibonaccichick666):

yes,but that still isnt 1

OpenStudy (anonymous):

My other problems is 1/1-sinx = 1 +sinx/ cos^2x

OpenStudy (fibonaccichick666):

that I can buy

OpenStudy (anonymous):

so 1 +sinx/ cos^2x will be 1+sinx/1-sin^2x, I need to get rid of the sinx/sinx

OpenStudy (fibonaccichick666):

do you mind posting it as another q though so I can get another medal?

OpenStudy (anonymous):

okay